OpenStudy (anonymous):

1. Riley noticed an aquarium in his father’s office. The aquarium had the dimensions 16 in. by 8.5 in. by 10.5 in. The formula for volume is: V = l × w × h. (a) Riley noticed that the aquarium was three-fourths full of water. How many more cubic inches of water would be required to fill the tank? Show your work. (b) Another aquarium in the building has dimensions that are each triple the size of the aquarium in Riley’s father’s office. Riley thought that the volume would also triple. Is Riley correct? How many times greater is the volume of the larger aquarium than the smaller?

OpenStudy (danjs):

1/4 of the volume is required to fill the tank (0.25) * (length x width x height)

OpenStudy (danjs):

If you triple each length in the tank V = 3*L x 3*W + 3*H = 3*3*3 (L X W X H) = 27 (L x W X H) = 27* Volume

OpenStudy (danjs):

In general, if you multiply the lengths by a constant number A, the volume will increase by \[A^3\] = A*A*A the original volume

OpenStudy (anonymous):

okay, so what your saying is find the volume first?

OpenStudy (danjs):

You can, if you want to show in numbers, for the second part. Find the volume, then find the volume using 3 times the lengths of each side

OpenStudy (danjs):

Volume of larger (3x) = 27*(Original Volume)

OpenStudy (anonymous):

81?

OpenStudy (danjs):

Original Volume = L*W*H Larger Volume = 3*L *3*W *3*H = (3*3*3)*(L*W*H) = 27*(Original Volume)

OpenStudy (anonymous):

729?

OpenStudy (danjs):

what did you do

OpenStudy (danjs):

Original Volume = 16 * 8.5 * 10.5 = 1428 in^3

OpenStudy (anonymous):

27 x 3 x 3 x 3

OpenStudy (danjs):

Larger Volume = 3*16 * 3*8.5 * 3*10.5 = (3*3*3) * (16*8.5*10.5) = 27* 1428 = 38556 in^3

OpenStudy (anonymous):

ok......how would i put that in a page that says show work?

OpenStudy (danjs):

\[\frac{ Larger ~Volume }{ Original~Volume } = \frac{ 38556 }{ 1428 } = 27\] The Larger volume is 27 times more than the original, if you triple the length of each side

OpenStudy (danjs):

Or you could just write Original Volume = Length x Width x Height Larger Volume = 3*Length x 3*Width x 3*Height = (3*3*3) *(Length x width x height) = 27* (Original Volume)

OpenStudy (anonymous):

thanks i have more

OpenStudy (anonymous):

2. Mr. Cooper’s company rents storage containers to customers. The customers can choose which size fits their needs and Cooper Storage delivers the containers. After the customers fill the containers, the company transports them to a warehouse to be stored until the customer needs the contents. (a) The volume for one size of the storage containers is 574 ft3. The container has a length of 10 ft and a width of 7 ft. How tall would a stack of 6 containers be if they were stacked one on top of another? (b) There are two other different sizes of containers. The price for renting a container per month for a residence within 30 miles is: 7’ container: 7 ft by 7 ft by 8 ft $99.00 12’ container: 12 ft by 8 ft by 8 ft $199.00 What is the price per cubic foot for each container? Round prices to the nearest cent.

OpenStudy (anonymous):

just tell me what to wright, and you will get a medal, since i have already fanned you @DanJS

OpenStudy (danjs):

start a new question please

OpenStudy (anonymous):

ok